Solve the equation: (3x-4)1/2=2


1
Expert's answer
2021-03-09T14:50:08-0500

"{(3x - 4)^{\frac{1}{2}}} = 2 \Rightarrow \left\{ {\begin{matrix}
{3x - 4 > 0}\\
{3x - 4 = {2^2}}
\end{matrix}} \right. \Rightarrow \left\{ {\begin{matrix}
{x > \frac{4}{3}}\\
{3x = 4 + 4}
\end{matrix}} \right. \Rightarrow \left\{ {\begin{matrix}
{x > \frac{4}{3}}\\
{x = \frac{8}{3}}
\end{matrix}} \right. \Rightarrow x = \frac{8}{3}"

Answer: "x = \frac{8}{3}"